# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ 4.4. Inductive estimates for $v_{q+1}$
In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ (4.16), it follows from (3.8d), (4.10) and (A.5) that
@@
where $M$ is the universal constant in (A.5). Taking expectation on the both side of (4.20), we then use (3.8e) to obtain
@@
where $\bar{M}$ is a universal constant satisfying $9c_{*}^{-\frac{1}{2}}M|\Lambda_{j}|+8M\bar{e}|\Lambda_{j}|<\bar{M}$ and we choose $a$ sufficiently large to absorb the constant in the above two inequalities. Applying (4.8b), (4.15) and (C.8), we also obtain
@@
where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant. We proceed to estimate $w^{(c)}_{q+1}$, using a standard mollification estimate, (4.8a) and (4.15) implies
@@
By (4.21) and (4.24), we have
$|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\frac{3}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$
[…]
$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$
Combining the above two estimates, we obtain
@@
4.5. Definition of the Reynolds stress $\mathring{R}_{q+1}$
[…]
$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.28)
$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes z_{q+1}-v_{q+1}\otimes z_{\ell}+z_{q+1}\otimes v_{q+1}-z_{\ell}\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
[…]
$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
In order to define the remaining oscillation error from the second line in (4.28), first note that for $j,j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$. Therefore, we may use (4.13) and Lemma A.2 to obtain
Note that compared to [BV19], a slight difference in this term is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$, which can be put into the pressure term $p_{\textrm{oscillation}}$. Thus, we have denoted
and
Finally we define the Reynolds stress at the level $q+1$ by
and
@@ 4.6. Inductive estimates for $\mathring{R}_{q+1}$
In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$. We first proceed with estimates of $R_{\mathrm{commutator1}}$ in subsection 4.6.1. Then we recall a useful sationary phase lemma in subsection 4.6.2. Finally, the remianing errors are controlled in subsection 4.6.34.6.6
@@ 4.6.1. Estimate on $R_{\mathrm{commutator1}}$
First, by the definition of $z_{q}$ we have
@@
where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds true for any $t\in\mathbb{R}$
@@
Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{com1}}$. Before proceeding, we give an estimate of $z_{q+1}-z_{q}$.
@@
By the definition of $z_{q}$, we deduce for any $p\geqslant 1$
@@
Finally, combining (4.35) with (4.36) implies
@@
In particular, taking $p=2r$ impies (4.33). Hence, the proof of Lemma 4.1 is completed. ∎
With Lemma 4.1 in hand, we proceed to estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$, using estimates (4.27), (4.34) and Hölder’s inequality, we have
@@
Using a standard mollification estimate, we obtain for any $\delta\in(0,\frac{1}{12})$
@@
Taking expectation on the both side and using (3.7), we deduce
@@
Hence, using (4.27), (4.37) and Hölder’s inequality, we deduce
@@
Similarly, using (3.7) and (4.37) again, we deduce
@@
Summarizing over the estimates above, we obtain
@@
$|\!|\!|\mathring{R}_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
where we used $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta-\frac{b}{8}\gamma\sigma}<\lambda_{q}^{-2b^{3}\beta}$. We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and we chose $a$ sufficiently large to absorb the constant.
@@ 4.6.2. Stationary phase Lemma
Before we estimate the remaining terms in (4.30) separately to show that the stress $\mathring{R}_{q+1}$ obeys the bound (3.8d), (3.8e) and (3.8f) at level $q+1$, we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$. We recall the following stationary phase lemma (see for example [BV19, Lemma 5.7] and [DS17, Lemma 2.2]), adapted to our setting.
@@
The above lemma is used when estimating the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.30). Indeed, for a fixed $t\in[k+(j-1)\ell,k+(j+1)\ell]$, in view of the bound (4.8b), we have $\frac{1}{2}\leqslant\|\Phi_{k,j}(\cdot,t)\|_{C^{0}_{x}}\leqslant 2$ on $\mathbb{T}^{3}$. Thus, it follows from Lemma 4.2 that if $a$ is a smooth periodic function, then we have for any $n\in\mathbb{N}$
@@
The implicit constant depends only on $\alpha$ and $n$. In the following subsection, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error. We first fix an integer $b\geqslant 7$, then choose $\beta$, $\alpha$ sufficiently small such that $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 and choose $m\in\mathbb{N}$ such that $m>\frac{2b+5}{b-5}$, at last we choose $a$ sufficiently large to absorb the constant.
@@
Inspecting the definition of $w_{q+1}^{(p)}$ as (4.16), we notice that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$. Therefore, for $t\in[k,k+1]$, $k\in\mathbb{Z}$ we have
@@
Returning to the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (4.14) and (4.15) and the definition of $z_{q}$ (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
@@
where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
Based on previous computations (4.29) in Section 4.5, we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
@@
Then using (4.14), (4.15) and chain rule, we have that for $n\in\mathbb{N}_{0}$
@@
where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$. Combining (4.43), (4.44) with (4.45) and using $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) again, applying (4.41) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$ implies
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$, and used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$, and we chose $a$ sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
Similar to (4.42), using (3.8a), (4.14), (4.15), (4.31) and chain rule, we have for $n\in\mathbb{N}_{0}$
@@
where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation, we obtain
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
The corrector error has two pieces, the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$, and the residual contribution from the nonlinear term, which are easier to estimate due to (4.20) and (4.23). Thus we obtain
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Combining the above estimates and taking expectation, we obtain
@@
The remaining term in corrector error is
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$, and used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Taking expetation and combining (4.50), we obatin
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
Summarizing over all the estimates we have established above and by (4.3), (4.4) and (4.32), we obtain

### v3

@@ 5.1. Inductive estimates for $v_{q+1}$
In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ in (4.11), it follows from (3.8d), (4.6) and (A.5) that
@@
where $M$ is the universal constant in (A.5). Taking expectation on both sides of (5.3), we use (3.8e) to obtain
@@
where $\bar{M}$ is a universal constant satisfying $9c_{*}^{-\frac{1}{2}}M|\Lambda_{j}|+8M\bar{e}|\Lambda_{j}|<\bar{M}$ and we choose $a$ sufficiently large to absorb the constant in the above two inequalities. Using (4.4b), (5.2) and (C.8), we obtain
@@
where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality. We also chose $a$ sufficiently large to absorb the constant.
We next estimate $w^{(c)}_{q+1}$. Using a standard mollification estimate, (4.4a) and (5.2) we obtain
@@
We use a standard mollification estimate and inductive assumption (3.8c) to obtain for any $t\in\mathbb{R}$
$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}},$ (5.9)
where we used $b\beta<1$ and chose $a$ sufficiently large to absorb the constant. Taking expectation we obtain
$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (5.10)
Combining (5.4), (5.7) and (5.10) we obtain
@@ 5.2. Inductive estimates for $\mathring{R}_{q+1}$
In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold at the level $q+1$. We first proceed with estimates of $R_{\mathrm{commutator1}}$ in subsection 5.2.1. Then we recall a stationary phase lemma in subsection 5.2.2. Finally, the remaining errors are controlled in subsection 5.2.35.2.6.
@@ 5.2.1. Estimate on $R_{\mathrm{commutator1}}$
First, by the definition of $z_{q}$ in (3.5), we have
@@
where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds for any $t\in\mathbb{R}$
@@
Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{commutator1}}$, and we first give an estimate of $z_{q+1}-z_{q}$.
@@
By the definition of $z_{q}$ in (3.5), we deduce for any $p\geqslant 1$
@@
where we used $9b\gamma\leqslant 8b^{2}\gamma+\gamma$ to have $\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\gamma/8}\leqslant\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\gamma/8}$. Finally, combining (5.17) with (5.18) implies
@@
In particular, taking $p=2r$ impies (5.15). Hence, the proof of Lemma 5.2 is complete. ∎
Next, we estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$. Using (5.12), (5.16) and Hölder’s inequality, we have
@@
We use a standard mollification estimate to have for any $\delta\in(0,\frac{1}{12})$
@@
Taking expectation on both sides and using (3.7), we deduce
@@
Hence, using (5.12), (5.19) and Hölder’s inequality, we have
@@
Similarly, we use (3.7) and (5.19) again to deduce
@@
Summarizing the estimates above, we obtain
@@
$|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
where we used $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta-\frac{b}{8}\gamma\sigma}<\lambda_{q}^{-2b^{3}\beta}$. We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and chose $a$ sufficiently large to absorb the constant.
@@ 5.2.2. Stationary phase Lemma
Before we estimate the remaining terms in (4.18), we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$. We recall the following stationary phase lemma (see for example [BV19, Lemma 5.7] and [DS17, Lemma 2.2]), adapted to our setting.
@@
The above lemma is used to estimate the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.18). Indeed, for a fixed $t\in[k+(j-1)\ell,k+(j+1)\ell]$, in view of the bound (4.4b), we have $\frac{1}{2}\leqslant\|\Phi_{k,j}(\cdot,t)\|_{C^{0}_{x}}\leqslant 2$ on $\mathbb{T}^{3}$. Thus, it follows from Lemma 5.3 that if $a$ is a smooth periodic function, then we have for any $n\in\mathbb{N}$
@@
The implicit constant depends only on $\alpha$ and $n$.
In the following subsections, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error. We first fix an integer $b\geqslant 7$, then choose $\beta$, $\alpha$ sufficiently small such that $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 and choose $m\in\mathbb{N}$ such that $m>\frac{2b+5}{b-5}$, at last we choose $a$ sufficiently large to absorb the constant.
@@
Inspecting the definition of $w_{q+1}^{(p)}$ as (4.11), we note that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$. Therefore, for $t\in[k,k+1]$, $k\in\mathbb{Z}$ we have
@@
Recalling the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (5.1) and (5.2) and the definition of $z_{q}$ in (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
@@
where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$. We also used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
Based on the previous computations in (4.17), we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
@@
Then using (5.1), (5.2) and chain rule, we have for any $n\in\mathbb{N}_{0}$
@@
where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely, $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$. Combining (5.25), (5.26), (5.27) with (C.7) again, applying (5.23) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$, we obtain
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$. We also used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$ and we chose $a$ sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
Using (3.8a), (5.1), (5.2), (5.13) and chain rule, we have for $n\in\mathbb{N}_{0}$
@@
where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$. We also used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$ and chose $a$ sufficiently large to absorb the constant. Then taking expectation, we obtain
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
@@
The corrector error has two pieces, one is the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$. The orther is a residual contribution from the nonlinear terms, which is easier to estimate due to (5.3) and (5.6). Therefore, we have
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$. Also, $a$ was chosen sufficiently large to absorb the constant. Combining the above estimates and taking expectations we obtain
@@
The remaining term in the corrector error is
@@
where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$. We also used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$ and chose $a$ sufficiently large to absorb the constant. Taking expetation and combining (5.32), we obatin
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},\qquad|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
5.2.7. Estimate on the commutator error
Using the definition of $z_{q}$ in (3.5) and the pathwise inductive assumption (3.8a), we obtain
$\|R_{\textrm{commutator}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (5.33)
where we chose $a$ sufficiently large to absorb the constant in the last inequality. Using a mollification estimate and inductive assumption (3.8a) and (3.8c), we have for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$
$\|R_{\textrm{commutator}}(t)\|_{L^{\infty}}$
Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
$|\!|\!|R_{\textrm{commutator}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (5.34)
$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. By (5.34), we also obtain
$\qquad|\!|\!|R_{\textrm{commutator}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (5.35)
Summarizing over all the estimates we have established above and by (5.14) and (5.33), we obtain